OpenStudy (anonymous):

MEDAL AND FAN. Let f(x)=–5x + 3 and g(x) = 6x – 2. Find f *g and its domain. I'm unsure of two things, do I FOIL? and how do I find the domain?

OpenStudy (anonymous):

@ganeshie8 do you mind??

OpenStudy (anonymous):

fxg =(-5x+3)(6x-2)=-30X^2+10X+18X-6= -30X^2+8X-6 DOMAIN=(-INFINITY , +INFINITY)

OpenStudy (anonymous):

Okay, so what do you mean by infinity?

OpenStudy (anonymous):

(-∞ , ∞ )

OpenStudy (anonymous):

\[(-\infty, +\infty)\]

OpenStudy (anonymous):

Okay.. but what do you mean by that?! I understand the multiplication but where does infinity come in?

OpenStudy (anonymous):

The domain is.

OpenStudy (anonymous):

@mikaylabean it means value of x can be from -infinty to + infinity. There is no specific domain

OpenStudy (anonymous):

How do you determine that? Because I have a list of possible answers, and of course the right one is "all real numbers" but.. What about when theres an answer? How did he get that?

OpenStudy (anonymous):

cuz that is a linear function and you know each linear function domain has a \[(-\infty, +)\]

OpenStudy (anonymous):

Okay, so then how do you determine if it is linear? I haven't reviewed that in two years.

OpenStudy (anonymous):

for instance, my next answer is "x-4" how would i determine that domain? Guys I'm really at a loss here.

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

Would you mind helping me with a harder one? @sangya21

OpenStudy (anonymous):

I can try.

OpenStudy (anonymous):

Okay. Let f(x)=x^2+2 and g(x)=x+8/x Find (g*f)(-7)

OpenStudy (anonymous):

Nothin?

OpenStudy (anonymous):

(g*f)(-7) = g(f(-7) \[g(f(x)) = \frac{ x^2 +16 +8 }{ x^2 +16 }\] Put x = -7 \[g(f(-7)) = \frac{ (-7)^2 +24 }{ (-7)^2 +16 }\]

OpenStudy (anonymous):

So the -7^2 would cancel right?

OpenStudy (anonymous):

nopes. its not product but addition \[g(f(-7) = \frac{ 49+24 }{ 49+16 }\] \[g(f(-7) = \frac{ 73 }{ 65 }\]

OpenStudy (anonymous):

Thats exactly what I got, but it's not one of the responses. The closest is 63/55

OpenStudy (anonymous):

what are the other options?

OpenStudy (anonymous):

\[-\frac{ 55 }{ 7 }, \frac{ 384 }{ 7 }, \frac{ 295}{ 49 }\]

OpenStudy (anonymous):

I did wrong \[g(f(x)) = \frac{ x^2 +2 + 8 }{ x^2 +2}\] \[g(f(-7)) = \frac{ (-7)^2 +2 + 8 }{ (-7)^2 +2}\] \[g(f(-7)) = \frac{ 59 }{ 51}\]

OpenStudy (anonymous):

This is getting very frustrating v.v

OpenStudy (anonymous):

It was 63/55

OpenStudy (anonymous):

Okkk

OpenStudy (anonymous):

sorry